males with less than 4 years of highschool,14 ,4years of highschool only 25,some college(less than four years20.4 years college or more23 total 82
Females in the same order are 15,31,24,22 ,total 92 and the total for each is 20,56,44,45,and total of 174 These were in a chart but dont know how to enter a chart.
Find the probability, expressed in a simplified fraction, that a randomly selected American, aged 25 or over
1. has not completed 4 yrs (or more) of college
2. has not completed 4 yrs of high school
3. has completed four years of high school only or less than 4 years of college
4. has completed less than 4 yrs of high school or four years of high school only
5. has completed 4 yrs of high school only or is a man
6. has completed 4 years of high school only or is a woman

Answer by ewatrrr(24785)


Hi
HS<4 HS=4 C<4 C≥4 Total
males 14 25 20 23 82
females 15 31 24 22 92
total 29 56 44 45 174
1. P(has not completed 4 yrs (or more) of college) = 129/174
2. P(has not completed 4 yrs of high school) = 29/174
3. P(has completed four years of high school only 'or' less than 4 years of college) 56/174 + 129/174 - 56/174
4. P(has completed less than 4 yrs of high school or four years of high school only)= 29/174 + 44/174 - 29/174
5. p(has completed 4 yrs of high school only 'or' is a man) 56/174 + 82/174 - 14/174.
6. P(has completed 4 years of high school only or is a woman)= 56/174 + 92/174-31/174